OpenStudy (anonymous):

In an election, 2.8 million votes were cast and each vote was for either Candidate I or Candidate II. Candidate I received 28,000 more votes than Candidate II. What percent of the 2.8 million votes were cast for Candidate I?

hartnn (hartnn):

where are u stuck ? could u find candidate I 's votes ?

OpenStudy (anonymous):

is the answer 51 percent?

hartnn (hartnn):

idk, i didn't calculate

OpenStudy (anonymous):

1,428,000/2,8000,000 equals .51=51 percent

hartnn (hartnn):

seems correct :)

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

its 50.5 percent

hartnn (hartnn):

huh ? how ?

hartnn (hartnn):

are u taking a quiz or something ?

OpenStudy (anonymous):

hold up

hartnn (hartnn):

ohh..x+x+28000 = 2.8 million yes, its 50.5%

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

so it set it up algebraical?

OpenStudy (anonymous):

2x+28000=2.8?

hartnn (hartnn):

2.8 million = 28 00000

hartnn (hartnn):

x= ?

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

dang

OpenStudy (anonymous):

I am not getting 50.5 percent...

hartnn (hartnn):

x= ?

OpenStudy (anonymous):

x=50

hartnn (hartnn):

!! x= 1386000

OpenStudy (anonymous):

wait...

OpenStudy (anonymous):

ahh... you are right

OpenStudy (anonymous):

so the second candidate received 49.5 percent of the votes

hartnn (hartnn):

1386000+28000=1414000 1414000/2.6mill*100=50.5

OpenStudy (anonymous):

and Candidate I received 50.5 percent of the votes

hartnn (hartnn):

yup

OpenStudy (anonymous):

next problem?

hartnn (hartnn):

yup